OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form.

jimthompson5910 (jim_thompson5910):

y = mx+b y = 2x+b 5 = 2(-1) + b 5 = -2 + b 5 +2 = b 7 = b b = 7 So the equation in slope intercept form is y = 2x + 7, now convert to standard form y = 2x + 7 y - 2x = 7 -2x+y = 7 2x - y = -7

OpenStudy (anonymous):

y-5=2(x+1) y=2x+2+5 y=2x+7 2x-y+7=0

OpenStudy (anonymous):